### v1

# Spatio-Temporal Adaptive Diffusion Models for EEG Super-Resolution in Epilepsy Diagnosis

Abstract
Electroencephalogram (EEG) technology, particularly high-density EEG (HD EEG) devices, is widely used in fields such as neuroscience. HD EEG devices improve the spatial resolution of EEG by placing more electrodes on the scalp, meeting the requirements of clinical diagnostic applications such as epilepsy focus localization. However, this technique faces challenges such as high acquisition costs and limited usage scenarios. In this paper, spatio-temporal adaptive diffusion models (STADMs) are proposed to pioneer the use of diffusion models for achieving spatial SR reconstruction from low-resolution (LR, 64 channels or fewer) EEG to high-resolution (HR, 256 channels) EEG. Specifically, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which then serve as conditional inputs to guide the reverse denoising process of diffusion models. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance to generate subject-adaptive SR EEG. Experimental results demonstrate that the proposed method effectively enhances the spatial resolution of LR EEG and quantitatively outperforms existing methods. Furthermore, STADMs demonstrate their value by applying synthetic SR EEG to classification and source localization tasks of epilepsy patients, indicating their potential to significantly improve the spatial resolution of LR EEG.
Spatio-temporal adaptive, diffusion models, super-resolution (SR), Transformer, source localization.
1 Introduction
Electroencephalogram (EEG) is a non-invasive, cost-effective neuroimaging technique with high temporal resolution, playing a crucial role in clinical diagnosis and cognitive neuroscience research. Unlike other neuroimaging techniques such as functional magnetic resonance imaging (fMRI) and positron emission tomography (PET), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability provides substantial support for studying dynamic cognitive processes [1]. In recent years, the emergence of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as sleep monitoring and brain-computer interface (BCI) control [2, 3, 4]. However, these devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [5]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [6, 7, 8]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [9]. These factors collectively lead to difficulties in acquisition and limited application scenarios for HD EEG devices.
The aforementioned observation highlights a significant dilemma in the current clinical diagnosis and treatment using EEG: due to high costs, discomfort from wearing, and other factors, high-resolution (HR) EEG is scarce. Consequently, the spatial resolution of low-resolution (LR) EEG is insufficient for clinical needs, particularly in the preoperative evaluation of patients with refractory epilepsy [10]. In such cases, the EEG spatial super-resolution (SR) method offers an effective remedy. This technique enables the reconstruction of HR EEG from LR EEG collected from subjects, thereby enhancing the spatial resolution of the existing EEG acquisition systems. For instance, EEG SR can reconstruct 256-channel EEG from 32-channel counterparts. Notably, EEG SR methods hold substantial potential for clinical applications, providing a cost-effective and efficient method to acquire HR EEG data.
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models, and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [11]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances [12] and multi-quadratic interpolation methods [13]. However, these methods only utilize the spatial correlations between adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity, usually yielding limited results for non-stationary signals.
In recent years, deep learning methods have become mainstream for analyzing and enhancing medical images [14, 15, 16]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [17], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [18] to correlate the structural and functional connections of participants’ brains [19]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [20, 21]. Existing researches have utilized approaches such as Autoencoders (AE) [22] and Generative Adversarial Networks (GANs) [23] to enhance the spatial resolution of EEG. While existing deep learning-based EEG SR methods have achieved significant improvements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. Moreover, the enhanced spatial resolution achieved by these methods is constrained, with a maximum capability of reconstructing EEG data up to 64-channel level, which may not meet the requirements of specific clinical scenarios.
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [24] have demonstrated formidable generative capabilities in the field of computer vision, sparking a surge in the Artificial Intelligence Generated Content (AIGC) [25, 26]. Moreover, diffusion models have been shown to surpass GANs in image generation quality, while also offering ideal characteristics such as distribution coverage, fixed training targets, and scalability [27]. Besides, Latent Diffusion Models (LDMs) [28] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. Moreover, SynDiff [29] and DiffMDD [30] have both demonstrated that diffusion models can be applied to the generation of medical images, thereby improving the accuracy of disease diagnosis. Despite this, the application of diffusion models in enhancing the spatial resolution of EEG remains relatively unexplored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications across various neuroscience contexts.
In this paper, a novel EEG SR method, called spatio-temporal adaptive diffusion models (STADMs), is proposed to achieve spatial SR reconstruction from LR EEG to HR EEG. This method enhances the spatial resolution of EEG by learning the latent mapping relationship between LR EEG and HR EEG through diffusion learning strategy. The main contributions of this paper are summarized as follows:
1.
A spatio-temporal adaptive framework based on diffusion models is proposed for EEG SR reconstruction. To the best of our knowledge, this is the first instance where diffusion models are employed to achieve spatial SR reconstruction from LR EEG to HR EEG. This provides an efficient and cost-effective method for obtaining HR EEG, enhancing the portability and usability of HR EEG, and further advancing fields such as epileptic focus localization and BCI technology.
2.
The spatio-temporal condition (STC) module is designed to extract the spatio-temporal features of LR EEG. It further uses the generated spatio-temporal conditions to constrain the reverse process of the diffusion model, guiding the model to generate SR EEG that conforms to the characteristics of the subjects.
3.
The multi-scale Transformer denoising (MTD) module is constructed to map LR EEG to HR EEG. It extracts temporal features at various scales and uses cross-attention-based diffusion Transformer blocks to adaptively modulate the denoising process based on spatio-temporal conditions, addressing the channel-level disparity between LR EEG and HR EEG.
2 Method
2.1 Preliminaries and Problem Statement
Before introducing our architecture, we briefly review some basic concepts needed to understand diffusion models [24]. Diffusion models assume a forward diffusion process which gradually applies noise to input data : , where are the variance schedule of the quantity of Gaussian noise.
In the reverse denoising process, the denoising model needs to learn the distributions to denoise the each state during training. Furthermore, the training of diffusion models is accomplished using a variational method-based training strategy [31]. Formally, the inference procedure can be described as a reverse Markov process that starts with Gaussian noise , and progresses towards the target data as:
Once is trained, new data can be sampled by initializing and sampling via the reparameterization trick, where denotes the number of forward process steps.
In this work, given an LR-HR EEG pair denoted as with being a degraded version of , where and denote the number of the LR and HR EEG channels, respectively. is the length of each EEG epoch. Therefore, in EEG SR tasks, the end-to-end mapping from the LR EEG to the corresponding SR EEG can be expressed as
where represent synthetic SR EEG. denotes the mapping function and represent the optimal parameters that are looked for to reconstruct the HR EEG.
2.2 Basic ideas
Learning the mapping relationship from LR EEG to HR EEG is challenging. To ensure that the generated results conform to the subject’s characteristics, the proposed STADMs use the spatio-temporal features of LR EEG as conditions and employs a cross-attention-based conditional injection mechanism to guide the denoising model for progressive denoising restoration. This mechanism embeds the spatio-temporal information of LR EEG into the high-dimensional latent space, indirectly guiding the model to learn the complex mapping relationship between LR EEG and HR EEG. This ensures that the synthetic SR EEG is highly consistent with HR EEG in both temporal and spatial domains, addressing the issue of the large channel-level disparity between LR EEG and HR EEG.
As Fig. 1 shows, the entire model framework consists of EEG pre-trained autoencoder, STC module and MTD module. During the training phase of STADMs, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors , and the state of each time step is gradually obtained through the forward process of the diffusion models. In the reverse process, the STC module first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTD module takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. The MTD module is trained by optimizing the mean squared error loss between the predicted noise and the actual noise, and finally generates SR EEG through progressive denoising. During the sampling phase of STADMs, the deterministic MTD module precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
2.3 Architectures
2.3.1 Pre-trained EEG Autoencoders
LDMs [28] have demonstrated that mapping input data into a latent space through a pre-trained encoder can significantly enhance the quality of high-resolution natural image generation and improve performance on other downstream tasks. Following the work of [32], we employ a Masked Autoencoder (MAE) for asymmetric latent space representation of HR EEG. Given that the input EEG consists of multi-channel time series signals, we utilize two-dimensional EEG data as input, where each row represents the time series data of a single channel, and each column represents the values of all channels at a specific time point. Initially, this module divides the EEG data of each channel into fixed-length windows and randomly masks them according to a specified percentage. The MAE then predicts the missing values based on the surrounding contextual cues. By reconstructing the masked signals, the pre-trained EEG encoder can learn the latent spatio-temporal representations of brain activity across different subjects. Detailed information on the architecture of the pre-trained EEG encoder can be found in [32].
2.3.2 Spatio-temporal Condition Module
In EEG signals, each channel represents the signals collected by electrodes placed at specific locations on the brain, recording electrical activity from different brain regions [33, 34]. The spatio-temporal condition module based on Transformer is constructed to capture the temporal correlations between time points and the spatial correlations between EEG channels. This module takes the time series from LR EEG and the spatial positions of the channels as inputs and outputs corresponding encoding vectors as conditional information for the denoising model, where represents the LR EEG. This module can extract the spatio-temporal features of LR EEG, providing richer and more precise contextual information for the denoising model.
As the yellow part of Fig. 1 shows, the STC module is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the channel spatial embedding layer encodes the channel spatial coordinates of LR EEG, outputting a set of spatial features. Then, we partition the input time series into different patches and feed them into the 1D convolutional block to preliminarily extract temporal features [35]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition . The 1D convolution block consists of a 1D convolution layer, a ReLU activation function, and batch normalization layer. The Transformer block consists of a multi-head self-attention (MSA) layer followed by feed-forward networks, both wrapped in residual connections and layer normalization.
2.3.3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like epileptic source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [36, 37], the MTD module employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks are introduced to capture multi-scale temporal features in the reverse denoising process.
As shown in the blue part of Fig. 1, the MTD module mainly consists of position encoding, multi-scale 1D convolution blocks and diffusion Transformer blocks. Each diffusion Transformer block includes a normalization layer, a MSA layer, a cross-attention layer and a feed-forward network. The MTD module takes , timestep and spatio-temporal condition as inputs, outputting the predicted noise at timestep . The multi-scale 1D convolution blocks extract temporal features at different scales from the latent vectors using 1D convolution layers with various kernel sizes and concatenate them along the feature dimension. The output of this convolution path can be represented as:
where Conv denotes performing a 1D convolution on each channel sequence and BN represents batch normalization. denotes the output of different 1D convoluntion layer and denotes the number of 1D convoluntion layers. Subsequently, each row of is treated as a token and input into the diffusion Transformer block. Through the MSA layer, information from different scales is integrated, and global spatio-temporal dependencies are captured. For the input , where is the sequence length and is the feature dimension. The output of MSA layer can be expressed as
where LN denotes the layer normalization operation. Furthermore, condition is incorporated into Transformer through a cross-attention layer. The output of cross-attention layer can be expressed as
where are projection matrices with learnable parameters. is the scaling factor. Finally, the output is processed by a feed-forward network to obtain the final fused feature representation, and a standard linear decoder is used to output a predicted noise. This enables the reparameterization trick to obtain the denoised latent for the previous time step. This mechanism effectively integrates the temporal and spatial features of EEG, providing rich context for subsequent denoising and reconstruction tasks, thereby skillfully achieving the mapping from LR EEG to HR EEG.
2.4 Loss Function
In the forward process of the diffusion model, we gradually add Gaussian noise to the clean input to generate a series of noisy data . In the reverse process, the traditional method is to train a denoising model to predict the state of the previous time step. Diffusion models draw on the principles of variation method [38], optimizing model parameters by maximizing a variational lower bound based on the evidence lower bound (ELBO):
where KL denotes the Kullback-Leibler divergence. However, directly predicting may lead to a complex and unstable training process. To simplify the training process, we use the denoising model to predict the noise at the current time step , and then minimize the mean squared error loss between the predicted noise and the ground-truth noise for each time step to learn the model parameters, where x denotes the input LR EEG. The final loss function can be expressed as:
where is the proposed multi-scale Transformer denoising model and represents the proposed STC module. Algorithm 1 displays the complete training procedure with the final loss function.
3 Experiment
3.1 Dataset and Preprocessing
In this study, we utilized a publicly available dataset (Localize-MI [39]) for model training and validation. The dataset consists of high-density EEG data collected from seven drug-resistant epilepsy patients (mean age=35.1, age=5.4), comprising a total of 61 sessions (average session duration per subject=8.71, session duration per subject=2.65). Additionally, the dataset includes spatial locations of stimulus contacts in MRI space, MNI152 space [40], and Freesurfer surface space [41], corresponding to each subject, as well as the digitized positions of 256 scalp EEG electrodes. We employed the MNE software for preprocessing of the raw data. Following the work of [39], we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) at a sampling rate of 8000 Hz for 256-channel EEG data, and deployed with the Geodesic Sensor Net using the HydroCel CleanLead electrode distribution system.
| Scaling Factor | Channels | Montage |
| 2 | 128 | EGI-128 montage |
| 4 | 64 | EGI-64 montage |
| 8 | 32 | EGI-32 montage |
| 16 | 16 | EGI-16 montage |
In the experiment, we downsampled the original high-density EEG data along the channel dimension to obtain corresponding low-resolution (LR) counterparts at different spatial resolutions, with scaling factors approximately set to 2, 4, 8, and 16. To simulate the acquisition of low-density EEG data in the real world, as Table 1 shows, we mapped each scaling factor to different numbers of channels in the EGI standard electrode systems[42], obtaining corresponding EEG electrode arrays. Thus, by using electrode systems with different channel ranks, we identified and removed the channels to be estimated, resulting in LR EEG data that match the preset channel levels. These LR-HR data pairs at different spatial resolutions were then created for model training.
3.2 Experiment Settings
3.2.1 Implement Details
The proposed STADMs are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. During model training, the Adam optimizer was employed with a batch size of 32, a learning rate of , and 300 training epochs. In the diffusion model framework, the forward process was set to 1000 time steps, with noise levels following sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with kernel sizes of 3, 5, 7, and 9, respectively. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
3.2.2 Metrics
To evaluate the quality of the synthetic EEG signals, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Mean Absolute Error (MAE), Normalized Mean Squared Error (NMSE), and Signal-to-Noise Ratio (SNR) [43]. These metrics provide comprehensive insights into the accuracy and reliability of the reconstructed signals compared to the real signals.
To quantitatively evaluate the performance improvement of SR EEG in downstream tasks, we used classification performance metrics including Accuracy (ACC), Precision, Recall, and F1 Score. Additionally, to assess the performance of SR EEG in source localization, we used Localization Error [7] as a quantitative measure of performance improvement in source localization.
3.3 Performance Evaluation on EEG SR Reconstruction
In this section, we evaluate the performance of the proposed STADMs for EEG spatial super-resolution reconstruction, primarily analyzing the generated results through various quantitative metrics and scalp topography maps.
3.3.1 Comparison With Other Existing EEG SR Methods
the proposed STADMs are compared with the existing EEG spatial SR methods to demonstrate the effectiveness of the synthetic HR EEG (256 channels). We mainly quantitatively analyze the reconstruction performance differences between the proposed STADMs and CNN-based methods (Deep-CNN [17]), autoencoder-based methods [Weighted gate layer autoencoders (WGLAE) [22], Deep autoencoder (DAE) [44]], and GANs-based method (EEGSR-GAN) [23].
Fig. 2 shows quantitative comparison results between different methods. As observed, the proposed STADMs achieve the best performance, significantly outperforming other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STADMs can maintain high signal quality and correlation. Moreover, STADMs significantly outperforms other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STADMs and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG spatial super-resolution methods, the proposed STADMs can effectively reconstruct high-quality high spatial resolution EEG (256 channels), achieving high spatial resolution enhancement of LR EEG.
(a) PCC
(b) MAE
(c) NMSE
(d) SNR
3.3.2 Comparison With Different Scale Factors
In order to evaluate the performance of the proposed STADMs in synthetic SR EEG signals from LR EEG under different scale factors, we employ four metrics for quantitative analysis of the model generative performance. Fig. 3 shows quantitative results for different scale factors. As the scale factor increases, PCC and SNR values decrease, indicating a weakening correlation between reconstructed SR EEG and ground truth; meanwhile, MAE and NMSE values increase, implying a growing difference. This suggests that STADMs’s reconstruction performance diminishes at larger scale factors due to limited channel information, making HR EEG reconstruction challenging. However, even at larger scale factors, STADMs’s performance remains comparable to smaller scale factors. For example, with a scale factor of 16, the PCC value is only 3.02 % lower than at a scale factor of 8, the MAE value differs by 1.68 %, and the SNR value by 3.19 %. This indicates STADMs can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STADMs effectively enhances LR EEG spatial resolution across different scale factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
(a) PCC
(b) MAE
(c) NMSE
(d) SNR
3.3.3 Qualitative results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands.We first use the Multitaper method to extract the power spectral density (PSD) of LR EEG and corresponding synthetic SR EEG in five frequency bands, including (0.5 4 Hz), (4 8 Hz), (8 13 Hz), (13 30 Hz), and (30 40 Hz or higher) [45]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 4 compares scalp topographies across different frequency bands between EEG at various channel levels and model-generated 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low (, , ) and high-frequency bands (, ), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. Overall, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, whereas LR EEG results in coarse active regions, making it challenging to pinpoint active brain areas.
3.4 Classification Evaluation on the Synthetic SR EEG
Enhancing the spatial resolution of EEG significantly improves the performance of EEG-based diagnostic systems. For epilepsy patients, HR EEG aids doctors in identifying abnormal waveforms, such as spikes and sharp waves during seizures. To evaluate the effectiveness of synthetic SR EEG in detecting abnormalities, we designed a binary classification experiment (normal vs. abnormal) using SR EEG. In the Localize-MI dataset, EEG signals during stimulation are labeled abnormal, and those before stimulation are labeled normal. We train the EEG-Net [46] separately on both LR EEG and the synthetic SR EEG, and conduct corresponding classification tests. This allows us to compare the performance of different spatial resolution LR EEG and synthetic SR EEG in the classification task.
Fig. 5 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STADMs outperforms LR EEG in all metrics, showing that STADMs enhances classification performance. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STADMs effectively enhance LR EEG spatial resolution, improving downstream classification tasks. Moreover, this provides a approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
| freq bands | EEG | Accuracy (AVGSTD %) | ||
| scale factor 4 | scale factor 8 | scale factor 16 | ||
| band | LR | 69.810.02 | 67.210.00 | 63.870.12 |
| SR | 74.420.21 | 71.830.33 | 69.420.13 | |
| band | LR | 71.120.04 | 69.780.12 | 64.150.00 |
| SR | 75.360.01 | 73.210.32 | 72.130.09 | |
| band | LR | 77.430.01 | 74.370.21 | 65.220.39 |
| SR | 81.250.04 | 79.540.19 | 71.310.04 | |
| band | LR | 78.230.11 | 77.890.02 | 73.240.23 |
| SR | 84.340.16 | 81.470.10 | 78.630.12 | |
| band | LR | 76.580.04 | 73.290.18 | 70.040.27 |
| SR | 80.150.12 | 78.190.07 | 76.140.17 | |
| all | LR | 83.920.24 | 81.590.24 | 77.850.15 |
| SR | 87.610.18 | 83.510.28 | 79.520.06 | |
We also analyze classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to n 5 features (n is the number of channels), and classification is based on these features. Table 2 shows classification results for LR and SR EEG across different frequency bands and scale factors. SR EEG outperforms LR EEG in all frequency bands. As the scale factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands (, , ) have worse performance than high-frequency bands (, ), but the average accuracy improvement is higher in low-frequency bands (5.06 %) compared to high-frequency bands (4.95 %). This indicates that the proposed method better preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
3.5 Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [47]. To evaluate the effectiveness of the proposed method in enhancing the spatial resolution of EEG and improving downstream task performance, we assess it using the source localization task, including both quantitative and qualitative analyses. For quantitative analysis, we use localization error [7] as the metric, quantified as the Euclidean distance between the source localization results and the true source positions, to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [48] as the source localization method, an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [49]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
Fig. 6 shows a comparison of localization error for different EEG types among all subjects in the dataset. The localization error for LR EEG is widely dispersed with higher values, indicating lower precision and greater variability. In contrast, SR EEG shows significantly lower and more concentrated localization errors, suggesting higher precision and stability. Although localization error of SR EEG is slightly higher than HR EEG, its accuracy is markedly improved, making it a viable substitute for HR EEG and reducing reliance on expensive high-density equipment. Therefore, the experimental results demonstrate that the synthetic SR EEG substantially enhances the spatial resolution and source localization accuracy of LR EEG.
Fig. 7 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results and blue circles show the results from the LR EEG. The LR EEG results exhibit significant deviations from the ground truth, whereas SR EEG results show minimal deviation, closely matching HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, the method not only enhances EEG spatial resolution but also supports high-precision EEG imaging analysis in clinical and research applications.
4 Discussion and Conclusion
In this work, STADMs are proposed to address the spatial resolution limitations of low-density EEG devices, particularly in clinical diagnostic applications, such as epilepsy focus localization. STADMs are the first to employ a diffusion model to achieve spatial SR reconstruction from LR EEG to HR EEG. To ensure that the generated results align with subject-specific characteristics, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which are then used as conditional inputs to guide the reverse denoising process. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance, bridging the gap between SR EEG and HR EEG. Qualitative and quantitative results indicate that STADMs effectively enhance the spatial resolution of EEG. Furthermore, classification and source localization experiments demonstrate their ability to improve EEG performance in practical scenarios.
Experimental results show that while the proposed STADMs outperform existing methods in reconstruction performance, larger scale factors reveal a significant difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 3. We identified several possible reasons for this: i) At larger scale factors, the available spatial information for the model is limited, making it difficult to accurately capture the spatial relationships between channels; ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model; iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of other physiological signals. In the future, we will further optimize this framework and apply it to practical clinical scenarios in epilepsy diagnosis, aiding doctors in the preoperative assessment for epilepsy focus resection surgery.